Amanda Reaves , a zookeeper, wants to estimate the average weight gain of elephants who are on a new diet plan. A preliminary study found the standard deviation was estimated to be 5.50 kilograms. Determine the sample size that is required to estimate a population mean to within 0.95 kilograms with a 90% confidence interval.
Answer #1

Solution :

Given that,

standard deviation =   =5.50

Margin of error = E = 0.95

At 90% confidence level

= 1 - 90%  

= 1 - 0.90 =0.10

/2 = 0.05

Z/2 = Z0.05 = 1.645

sample size = n = [Z/2* / E] 2

n = ( 1.645 *5.50 / 0.95 )2

n =90.7

Sample size = n =91